# Transfer Learning for Transient Classification: From Simulations to Real Data and ZTF to LSST

### v1

@@ 4.2 ZTF to LSST
Figure 3: Classification performance as a function of training set size for models trained on LSST simulations. The orange points show models trained directly on LSST Sims, while the blue points shows models first trained on ZTF Sims and then fine-tuned on LSST sims. For transfer learning, we unfreeze the initial GRU layers to handle the additional LSST passbands while keeping the dense layers frozen. Transfer learning achieves 95% of baseline performance while requiring only 30% of the training data. Error bars show the standard deviation across 10 independent training runs. The increased variance compared to Fig. 2 reflects the additional complexity of adapting to LSST’s six passbands. We use the Area Under the Receiver Operating Characteristics Curve (AUROC) to measure classification performance.
To evaluate transfer learning between surveys, we compare models pretrained on ZTF Sims with those trained directly on LSST simulations. Unlike the BTS experiment where we maintained real-world class distributions, here we sample an equal number of objects per class. This approach is possible with simulated data where we have abundant examples of even rare transient classes, and it enables us to evaluate performance across all classes including rare ones.
As shown in Fig. 3, transfer learning significantly reduces the data requirements for developing LSST classifiers. A model pretrained on ZTF achieves 95% of the baseline performance while requiring only 30% of the training data. Despite the additional complexity of adapting to LSST’s six passbands (compared to ZTF’s two), the pretrained model effectively leverages its understanding of transient behaviour to reduce the amount of new training data needed. This suggests that existing classification models from current surveys like ZTF can be efficiently adapted for LSST, enabling rapid deployment of reliable classifiers early in the survey.
@@ 4.3 Training Time and Computational Efficiency
Another advantage of transfer learning beyond reducing data requirements is reduced training time, which is especially important in the era of data-driven astronomy. We find that models initialized with pretrained weights converge 25% faster than those trained from scratch for two reasons. First, the model begins with an understanding of general transient behaviour, requiring fewer iterations to learn class-specific features. Second, with many layers frozen, there are fewer parameters to optimize during training.
@@ 5 Conclusion
    Fine-tuning existing models requires 70% to 95% less labelled data than training from scratch. A model pretrained on ZTF simulations achieves equivalent performance on real data while requiring only 25% of the labelled examples. Similarly, when adapting ZTF models for LSST, transfer learning maintains 95% of the baseline performance with just 30% of the training data. These results suggest that classifiers trained on existing surveys or simulations can be efficiently adapted to new data sources.
@@
The success of transfer learning between both simulated and real data, as well as between different surveys, demonstrates that these models learn generalizable features of astronomical transients. This has important implications for future surveys like LSST, suggesting that reliable automated classification will be possible soon after observations begin, rather than waiting to accumulate large training sets.
Several promising directions remain for future work. First, combining data from multiple existing surveys could provide more robust base models that capture a wider range of transient behaviour. Second, systematic investigation of different transfer learning architectures (e.g. which layers to freeze, optimal MLP configurations) could further improve performance. Finally, applying these techniques to a broader range of astronomical phenomena could extend the benefits of transfer learning beyond transient classification.
@@ Acknowledgements
We thank Nabeel Rehemtulla and Ved Shah for useful discussions and for providing access to the Bright Transient Survey (BTS) and ELAsTiCC datasets, respectively.
@@
The ZTF simulations were generated using models from PLAsTiCC (Kessler et al., 2019b), available at https://zenodo.org/record/2612896#.YYAz1NbMJhE. The light curves were simulated to match ZTF observing properties using the SNANA software package (Kessler et al., 2009) with ZTF survey observing logs. These simulations build upon those used in Muthukrishna et al. (2019), updated to resolve issues with core-collapse supernovae as described in Muthukrishna et al. (2021) and Gupta et al. (2024, 2025). The simulated data is available upon reasonable request to the corresponding author.
The ELAsTiCC simulations were provided to the authors by Ved Shah and are also publicly available from the LSST Dark Energy Science Collaboration (DESC) on the National Energy Research Scientific Computing Center (NERSC) at https://portal.nersc.gov/cfs/lsst/DESC_TD_PUBLIC/ELASTICC/.
The ZTF Bright Transient Survey (BTS) dataset was provided to the authors by Nabeel Rehemtulla and is also publicly available from the LSST alert brokers as described at https://sites.astro.caltech.edu/ztf/bts/bts.php.

### v2

@@ 4.2 ZTF to LSST
To evaluate transfer learning between surveys, we compare models pretrained on ZTF Sims with those trained directly on LSST simulations.
As depicted in Figure 2 [right], transfer learning significantly reduces the data requirements for developing LSST classifiers. A model pretrained on ZTF achieves 94% of the baseline performance while requiring only 30% of the training data. Despite the additional complexity of adapting to LSST’s six passbands (compared to ZTF’s two), pretraining helps the model learn features and capture some survey-agnostic patterns that are intrinsic to the astrophysical phenomena, thus reducing the amount of new training data required. This suggests that existing classification models from current surveys such as ZTF can be efficiently adapted for LSST, enabling the rapid deployment of reliable classifiers early in the survey.
@@ 4.3 Training Time and Computational Efficiency
Another advantage of transfer learning beyond reducing data requirements is reduced training time, which is especially important in the era of data-driven astronomy. We find that models initialised with pretrained weights converge 25% faster than those trained from scratch for two reasons. First, the model begins with an understanding of general transient behaviour, requiring fewer iterations to learn class-specific features. Second, with many layers frozen, there are fewer parameters to optimise during training.
@@ 5 Conclusion
    Fine-tuning existing models requires 70% to 95% less labelled data than training from scratch. A model pretrained on ZTF simulations achieves equivalent performance on real data while requiring only 5% of the labelled examples. Similarly, when adapting ZTF models for LSST, transfer learning maintains 94% of the baseline performance with just 30% of the training data. These results suggest that classifiers trained on existing surveys or simulations can be efficiently adapted to new data sources.
@@
The success of transfer learning between both simulated and real data, as well as between different surveys, demonstrates that these models learn generalisable features of astronomical transients. This has important implications for future surveys like LSST, suggesting that reliable automated classification will be possible soon after observations begin, rather than waiting to accumulate large training sets.
Based on the expected observation rates from the ELAsTiCC dataset, which simulates the first three years of LSST operations, the survey is expected to observe approximately 50,000 transient events in its first month. Since our results demonstrate that effective classifier performance requires as few as 3,000 labelled transients, we advocate for aggressive spectroscopic follow-up efforts to enable reliable automated classification within the first few weeks of LSST operations.
Several promising directions remain for future work. First, combining data from multiple existing surveys could provide more robust pretrained models that capture a wider range of transient behaviour. Furthermore, applying these techniques to a broader range of astronomical phenomena could extend the benefits of transfer learning beyond transient classification.
@@ Acknowledgements
Zwicky Transient Facility access for N.R. and V.S. was supported by Northwestern University and the Center for Interdisciplinary Exploration and Research in Astrophysics (CIERA). N.R. is supported by DoE award #DE-SC0025599.
@@
The ZTF simulations were generated using models from PLAsTiCC (Kessler et al., 2019), available at https://zenodo.org/record/2612896#.YYAz1NbMJhE. The light curves were simulated to match ZTF observing properties using the SNANA software package (Kessler et al., 2009) with ZTF survey observing logs. These simulations build upon those used in Muthukrishna et al. (2019), updated to resolve issues with core-collapse supernovae as described in Muthukrishna et al. (2021) and Gupta et al. (2024, 2025). The simulated data is available upon reasonable request to the corresponding author.
The ELAsTiCC simulations are publicly available from the LSST Dark Energy Science Collaboration (DESC) on the National Energy Research Scientific Computing Center (NERSC) at https://portal.nersc.gov/cfs/lsst/DESC_TD_PUBLIC/ELASTICC/.
The BTS source list is publicly available on the BTS sample explorer (https://sites.astro.caltech.edu/ztf/bts/bts.php) and light curves can be obtained from ZTF alert brokers.
Appendix A Class-based Performance
A.1 BTS
Figure 3: Class-specific classification AUROCs for transfer learning from simulated ZTF data to real BTS data.
Figure 3 shows the class-specific performance of transfer learning compared to direct training on BTS data. Transfer learning provides significant improvement across all transient classes, with the most pronounced benefits observed for rare classes such as SLSNe-I and TDEs, which have very few training samples.
A.2 LSST
Figure 4: Class-specific classification AUROCs for transfer learning from ZTF sims to LSST sims.
Figure 4 shows the class-specific performance for transfer learning from ZTF simulations to LSST simulations. The relative performance across different transient classes reflects trends observed in other photometric classifiers such as RAPID (Muthukrishna et al., 2019), ATAT (Cabrera-Vives et al., 2024), and ORACLE (Shah et al., 2025). Classes with distinctive photometric signatures, such as AGNs with their characteristic stochastic variability over long timescales, achieve higher classification accuracy than supernova subtypes with more subtle differences. For example, distinguishing between SN-Ib and SN-Ic based solely on photometric rise times remains challenging, as these subtle features do not always manifest clearly in broad-band observations.
Appendix B Performance without Minority Classes
Figure 5: A recreation of Figure 2 [left] with TDEs and SLSNe-I omitted.
The transfer learning performance improvement shown in Figure 2 [left] is partly attributed to minority classes (SLSN-I and TDE), which have very few labelled observations but are relatively easy to classify (Muthukrishna et al., 2019; Gupta et al., 2025). To test whether transfer learning benefits extend beyond these rare classes, Figure 5 shows the same analysis using only the three most common transient types (SNIa, SNII, and SNIb/c). Transfer learning still provides substantial improvement even when focusing solely on these well-represented classes.